OpenStudy (anonymous):

16xy^2/27z^5 all under a square root

OpenStudy (anonymous):

x=sqrt(x^2) sqrt(16xy^2/27z^5) =(sqrt(4sqrt(x)y /(sqrt(27) z^(5/2))^2) =4sqrt(x)y /(sqrt(27) z^(5/2)

OpenStudy (anonymous):